# PLUM: Improving Code LMs with Execution-Guided On-Policy Preference Learning Driven By Synthetic Test Cases

### v3

@@ A.3 Additional Results: What if directly using competitive coding datasets?
There are existing competitive coding datasets like Code Contests Li et al. (2022) which are equipped with test cases. Using more general datasets like OSS-Instruct (Wei et al., 2023), paired with synthetic test cases, is more advantageous for preference learning in code language models than using competitive coding datasets like Code Contests. Competitive coding datasets present complex problems with intricate edge cases, which can overwhelm the model and obscure fundamental instruction-following and preference-learning goals. In contrast, OSS-Instruct provides more accessible, more uniform instructions that allow for cleaner and more straightforward alignment. This helps models learn functional correctness more effectively without being distracted by the nuances of competitive coding. Additionally, OSS-Instruct, sourced from real-world open-source projects, avoids domain-specific biases that can arise from competitive coding, making it more generalizable and applicable across diverse programming environments.
@@ A.5 Results On Base Models
Model Type MBPP MBPP+ HE HE+ Avg. (Base) Avg. (+) Avg. (All) StarCoder2-Base Baseline 54.4 45.6 35.4 29.9 44.9 37.8 41.3 SFT 62.2 49.4 41.5 35.4 51.9 50.6 47.1 PLUM-KTO 60.4 49.1 46.3 39.6 53.4 51.2 62.2 CodeQwen-Base Baseline 72.2 60.2 51.8 45.7 62.0 53.0 57.5 SFT 73.4 62.4 67.7 59.1 70.6 66.5 65.7 PLUM-KTO 75.4 62.9 70.1 62.2 72.8 67.8 67.7 DS-Coder-Base Baseline 70.2 56.6 47.6 39.6 58.9 57.8 53.5 SFT 71.7 57.1 56.1 48.8 63.9 60.5 58.4 PLUM-KTO 72.9 58.9 56.7 48.8 64.8 61.9 59.3

### v4

@@ A.3 Additional Results: What if directly using competitive coding datasets?
There are existing competitive coding datasets like Code Contests [26] which are equipped with test cases. Using more general datasets like OSS-Instruct [49], paired with synthetic test cases, is more advantageous for preference learning in code language models than using competitive coding datasets like Code Contests. Competitive coding datasets present complex problems with intricate edge cases, which can overwhelm the model and obscure fundamental instruction-following and preference-learning goals. In contrast, OSS-Instruct provides more accessible, more uniform instructions that allow for cleaner and more straightforward alignment. This helps models learn functional correctness more effectively without being distracted by the nuances of competitive coding. Additionally, OSS-Instruct, sourced from real-world open-source projects, avoids domain-specific biases that can arise from competitive coding, making it more generalizable and applicable across diverse programming environments.
@@ A.5 Results On Base Models
Model Type MBPP MBPP+ HE HE+ Avg. (Base) Avg. (+) Avg. (All)
StarCoder2-Base Baseline 54.4 45.6 35.4 29.9 44.9 37.8 41.3
SFT 62.2 49.4 41.5 35.4 51.9 50.6 47.1
PLUM-KTO 60.4 49.1 46.3 39.6 53.4 51.2 62.2
CodeQwen-Base Baseline 72.2 60.2 51.8 45.7 62.0 53.0 57.5
SFT 73.4 62.4 67.7 59.1 70.6 66.5 65.7
PLUM-KTO 75.4 62.9 70.1 62.2 72.8 67.8 67.7
DS-Coder-Base Baseline 70.2 56.6 47.6 39.6 58.9 57.8 53.5
SFT 71.7 57.1 56.1 48.8 63.9 60.5 58.4
PLUM-KTO 72.9 58.9 56.7 48.8 64.8 61.9 59.3